# The Coverage Principle: A Framework for Understanding Compositional Generalization

### v1

@@ E.2 Step 2: probability of observing $k$ evidences for one fixed pair
$\Pr[Y_{ij}\geq k]\leq\Pr[Y_{ij}^{\star}\geq k]=\sum_{r=k}^{\infty}\frac{e^{-\mu}\mu^{r}}{r!}\leq\frac{\mu^{k}}{k!}\bigl{(}1+o(1)\bigr{)}.$
This upper bound $p_{k}:=\mu^{k}/k!(1+o(1))$ will be the edge probability used in the connectivity threshold of Step 3.
E.3 Step 3: connectivity inside each equivalence class
With the edge probability $p_{k}$ from Step 2, we now analyze when the $k$-evidence graphs become connected. Recall that under our learner assumption, perfect generalization requires every equivalence class to form a connected component in the $k$-evidence graph. We model this as a random graph connectivity problem.
Random graph construction
Fix one balanced class $E_{b}$ of size $n:=|{\mathcal{X}}|$ (Assumption (A1)). Create a graph $G_{b}$ whose vertices are the first-hop pairs in $E_{b}$, and place an undirected edge $\{i,j\}$ exactly when the pair $(i,j)$ has been observed in at least $k$ distinct shared contexts. By Step E.2, each potential edge appears with probability
$p_{k}=\Pr\bigl{[}Y_{ij}\geq k\bigr{]}=\frac{\mu^{k}}{k!}\,,\qquad\mu=\frac{N^{2}}{|{\mathcal{X}}|^{5}}.$
Why $G_{b}$ can be approximated as Erdős–Rényi
As mentioned earlier, the dependence among distinct edges in $G_{b}$ arises from the constraint that the total sample size is $N$, which induces a negative correlation. Such negative dependence reduces the likelihood of simultaneously creating many edges. Consequently, viewing $G_{b}$ as an independent Erdős–Rényi graph $G(n,p_{k})$ provides a conservative model, and any threshold we derive for connectivity under independence remains valid (or becomes easier to satisfy).
Classwise connectivity threshold
For Erdős–Rényi graphs, the classical result of Erdős–Rényi [97] states
$\Pr\bigl{[}G(n,p)\text{ is connected}\bigr{]}\xrightarrow[n\to\infty]{}1\quad\Longleftrightarrow\quad p\geq\frac{\log n+\omega(1)}{n}.$
Setting $n=|{\mathcal{X}}|$ and $p=p_{k}$ yields the requirement
$p_{k}=\frac{\mu^{k}}{k!}\geq\frac{\log|{\mathcal{X}}|}{|{\mathcal{X}}|}(1+o(1)).$
Substituting $\mu=N^{2}/|{\mathcal{X}}|^{5}$ and rearranging gives
$\frac{N^{2k}}{|{\mathcal{X}}|^{5k}\,k!}\gtrsim\frac{\log|{\mathcal{X}}|}{|{\mathcal{X}}|},\qquad\Longrightarrow\qquad N\gtrsim|{\mathcal{X}}|^{\frac{5k-1}{2k}}\,\bigl{(}k!\log|{\mathcal{X}}|\bigr{)}^{1/(2k)}.$
Because $(k!\log|{\mathcal{X}}|)^{1/(2k)}=(\log|{\mathcal{X}}|)^{O(1/k)}$ grows only poly-logarithmically, we hide it inside a $\tilde{\Omega}(\,\cdot\,)$:
@@ E.3 Step 3: connectivity inside each equivalence class
Since every class must be connected to achieve the data–sufficiency criterion in the problem setting, we conclude

### v2

@@ F.2 Main Theorem: Sample Complexity Bound for Perfect Coverage With High Probability
In particular, these two theorems can be easily combined (hence the detailed proof is skipped) in the regime of almost balanced cardinalities as $\Theta(n)$, which is described in the following corollary.
Corollary F.9 (Power-Law Sample Complexity Upper Bound, Simple).
Under the two-hop task setup described in Sec.˜F.1, Assume that ${\mathcal{X}}_{1}$, ${\mathcal{X}}_{2}$, ${\mathcal{X}}_{3}$, and $V_{b}$ ($\forall b\in B$) are all the sets of size $\Theta(n)$. Fix any $k\geq 1$ and any $\delta>0$ Then, there exists $N_{\delta}(n)>0$ satisfying $N_{\delta}(n)\leq O\left(\max\left\{n^{2.5-\frac{0.5}{k}}\cdot(k!\cdot\ln n)^{\gamma},\;\ln\frac{2}{\delta}\right\}\right),\qquad\text{with }\gamma=\begin{dcases}1,&\text{for }k=1;\\ \frac{1}{2k},&\text{for }k\geq 2,\end{dcases}$ such that, for any large enough $n$ and any $N>N_{\delta}(n)$, the learner with a uniformly randomly sampled training dataset $D$ (with replacements) achieves a perfect coverage of in-domain data, i.e., $\overline{D}\subset\mathrm{Cover}_{k}(D)$, with probability at least $1-\delta$.
Now, we prove our main theorem for $k\geq 2$ (Theorem˜F.7). After that, we will present the proof of the other case ($k=1$, Theorem˜F.8), which follows a similar plot except for a few steps at the end.
Proof of Theorem˜F.7.
We begin the proof with the following observation: a sufficient condition for a perfect coverage is that every $b$-evidence graph (Def.˜F.5) is a connected graph.
Lemma F.10 ($\downarrow$).
If all ${\mathcal{G}}_{b}^{(D,k)}$ are connected graphs ($\forall b\in B_{D}$), then we have $\overline{D}\subset\mathrm{Cover}_{k}(D)$.
Refer to Sec.˜F.3 for the proof. Thanks to the lemma above, it suffices to show the following with sufficiently large $N=\left|D\right|$:
@@
To this end, consider the multinomial random vector $(Z_{\bm{x}})_{{\bm{x}}\in{\mathbb{X}}}$ corresponding to a randomly sampled training dataset $D$, as defined in Proposition˜F.1. Define a property $\mathcal{C}$} as
$(Z_{\bm{x}})_{{\bm{x}}\in{\mathbb{X}}}\in{\mathcal{C}}\iff{\mathcal{G}}_{b}^{(D,k)}\text{ is connected }\forall b\in B_{D}.$ (7)
Then, we want to compute an upper bound of a probability ${\mathbb{P}}_{N}\left((Z_{{\bm{x}}})_{{\bm{x}}\in{\mathbb{X}}}\notin{\mathcal{C}}\right)$. Observe that the property $\mathcal{C}$} is a non-decreasing property (Def.˜F.11): if a vector $(Z_{\bm{x}})_{{\bm{x}}\in{\mathbb{X}}}$ satisfies $\mathcal{C}$}, then it still satisfies $\mathcal{C}$} after increasing in its entries by non-negative amounts. In other words, the negation of $\mathcal{C}$} (“not satisfying $\mathcal{C}$}”) is a non-increasing property. This is because adding a data point to the training dataset $D$ does not remove any edge from all evidence graphs, thereby preserving the connectivity. In general, we define the monotone property of vectors as below.
Definition F.11 (Monotone property).
We refer to a property $\mathcal{A}$} of $m$-dimensional vectors as a non-decreasing property when, for any vector $(\Delta_{1},\cdots,\Delta_{m})$ with non-negative entries $\Delta_{i}\geq 0$ ($\forall i=1,\cdots,m$),
$(z_{1},\cdots,z_{m})\in{\mathcal{A}}\implies(z_{1}+\Delta_{1},\cdots,z_{m}+\Delta_{m})\in{\mathcal{A}}.$ (Non-Decreasing Property)
On the other hand, we define non-increasing property $\mathcal{A}$} as what satisfies
$(z_{1},\cdots,z_{m})\in{\mathcal{A}}\implies(z_{1}-\Delta_{1},\cdots,z_{m}-\Delta_{m})\in{\mathcal{A}}.$ (Non-Increasing Property)
Taking advantage of the monotonicity, we can apply the Poissonization technique (Sec.˜F.3.1) to obtain an upper bound on the probability ${\mathbb{P}}_{N}\left((Z_{{\bm{x}}})_{{\bm{x}}\in{\mathbb{X}}}\notin{\mathcal{C}}\right)$. The Poissonization technique we use is summarized as the lemma below, which will be proved in Sec.˜F.3.1:
Lemma F.12 (De-Poissonizaiton of Monotone Multinomial Events, $\downarrow$).
Fix any $n\geq 1$. Define
$P_{n}:={\mathbb{P}}_{(Z_{1},\cdots,Z_{m})\sim\mathrm{Multinomial}(n;\,p_{1},\cdots,p_{m})}\left(\left(Z_{1},\cdots,Z_{m}\right)\in{\mathcal{A}}\right).$
Let $\mathcal{A}$} be a non-decreasing property (Def.˜F.11) of $m$-dimensional vector. Then, we have an upper bound and a lower bound for $P_{n}$ as follows:
$P_{n}$ $\leq{\mathbb{P}}_{\mathrm{Po}(n+{\varepsilon})}\left(\left(Z_{1},\cdots,Z_{m}\right)\in{\mathcal{A}}\right)+\exp\left(\textstyle-\frac{(3-c){\varepsilon}^{2}}{6n}\right),$ $(\forall c\in(0,3),~\forall{\varepsilon}\in(0,cn))$
$P_{n}$ $\geq{\mathbb{P}}_{\mathrm{Po}(n-{\varepsilon})}\left(\left(Z_{1},\cdots,Z_{m}\right)\in{\mathcal{A}}\right)-\exp\left(\textstyle-\frac{{\varepsilon}^{2}}{2n}\right).$ $(\forall{\varepsilon}\in(0,n))$
If $\mathcal{A}$} is a non-increasing property (Def.˜F.11), then we have similar upper and lower bounds for $P_{n}$:
$P_{n}$ $\leq{\mathbb{P}}_{\mathrm{Po}(n-{\varepsilon})}\left(\left(Z_{1},\cdots,Z_{m}\right)\in{\mathcal{A}}\right)+\exp\left(\textstyle-\frac{{\varepsilon}^{2}}{2n}\right),$ $(\forall{\varepsilon}\in(0,n))$
$P_{n}$ $\geq{\mathbb{P}}_{\mathrm{Po}(n+{\varepsilon})}\left(\left(Z_{1},\cdots,Z_{m}\right)\in{\mathcal{A}}\right)-\exp\left(\textstyle-\frac{(3-c){\varepsilon}^{2}}{6n}\right).$ $(\forall c\in(0,3),~\forall{\varepsilon}\in(0,cn))$
Here, we denote by ${\mathbb{P}}_{\mathrm{Po}(\lambda)}$ the probability measure under $Z_{i}\overset{\text{\emph{indep.}}}{\sim}\mathrm{Poisson}(\lambda p_{i})$ ($\forall i=1,\cdots,m$).
In particular, considering the disconnectedness ($\exists b\in B_{D}$) as a non-increasing property, we have
$\ {\mathbb{P}}_{N}\left(\exists b\in B_{D}\text{ such that }{\mathcal{G}}_{b}^{(D,k)}\text{ is disconnected}\right)$
$\leq{\mathbb{P}}_{\mathrm{Po}(N-{\varepsilon})}\left(\exists b\in B_{D}\text{ such that }{\mathcal{G}}_{b}^{(D,k)}\text{ is disconnected}\right)+\exp\left(-\frac{{\varepsilon}^{2}}{2N}\right)$
$\leq\sum_{b\in B_{D}}{\mathbb{P}}_{\mathrm{Po}(N-{\varepsilon})}\left({\mathcal{G}}_{b}^{(D,k)}\text{ is disconnected}\right)+\exp\left(-\frac{{\varepsilon}^{2}}{2N}\right)$ $(\because\text{union bound})$
for any positive number ${\varepsilon}<N$, where ${\mathbb{P}}_{\mathrm{Po}(N-{\varepsilon})}$ is the probability measure under i.i.d. Poisson random variables $Z_{\bm{x}}\sim\mathrm{Poisson}\left(\frac{N-{\varepsilon}}{\left|{\mathbb{X}}\right|}\right)$ ($\forall{\bm{x}}\in{\mathbb{X}}$). Let us take ${\varepsilon}=\sqrt{2N\ln\left(\frac{2}{\delta}\right)}$, which is smaller than $N$ by our choice of $N_{\delta}$ and ensures that
$\exp\left(-\frac{{\varepsilon}^{2}}{2N}\right)=\exp\left(-\ln\left(\frac{2}{\delta}\right)\right)=\frac{\delta}{2}.$
Now, we claim that each $b$-evidence graph ${\mathcal{G}}_{b}^{(D,k)}$ is a binomial random $k$-intersection graph (Sec.˜F.3.2) for every $b\in B_{D}$, under the Poissonization governed by ${\mathbb{P}}_{\mathrm{Po}(N-{\varepsilon})}$.
Lemma F.13 (Evidence Graphs are Binomial $k$-Intersection Graphs under Poissonization, $\downarrow$).
Let any $b\in B_{D}$. Consider a vector $(Z_{\bm{x}})_{{\bm{x}}\in{\mathbb{X}}}$ associated with a dataset $D$ (i.e., an input sequence $\bm{x}$} is sampled $Z_{\bm{x}}$ times in $D$). Let ${\mathbb{P}}_{\mathrm{Po}(\lambda)}$ be a probability measure for $Z_{\bm{x}}\overset{\text{i.i.d.}}{\sim}\mathrm{Poisson}\left(\lambda/\left|{\mathbb{X}}\right|\right)$. Then, under ${\mathbb{P}}_{\mathrm{Po}(\lambda)}$, the $b$-evidence graph ${\mathcal{G}}_{b}^{(D,k)}$ (Def.˜F.5) is an instance of binomial random $k$-intersection graph ${\mathcal{G}}^{(k)}(n_{b},m,p)$ with parameters
$n_{b}=\left|V_{b}\right|,\quad m=\left|{\mathcal{X}}_{3}\right|,\quad p=1-\exp\left(-\frac{\lambda}{\left|{\mathbb{X}}\right|}\right).$
Based on such an observation, we can use the following seminal result about the connectivity of binomial random $k$-intersection graphs with $p=1-\exp(-(N-{\varepsilon})/\left|{\mathbb{X}}\right|)$. In particular, we will use only the $k\geq 2$” part (which tightly matches our assumption $m=\Omega(n_{b})$ ($\forall b\in B_{D}$)) below:
Lemma F.14 (zhao2014k, Theorem 2; zhao2017connectivity, Theorem 1 & Remark 1).
Fix any $k\geq 1$. Suppose that
$m=\begin{cases}\Omega\left(\min\left\{n\left(\ln n\right)^{5},n^{\rho}\right\}\right),&\text{if }k=1,\text{ for any }\rho>1;\\ \Omega\left(n\right),&\text{if }k\geq 2,\end{cases}$
and
$p=\left(\frac{k!\,\left(\ln n+\alpha_{n}\right)}{n}\right)^{\frac{1}{2k}}\cdot\frac{1}{\sqrt{m}}$ (8)
for any sequence $\{\alpha_{n}\}$ which attains a limit $\alpha_{\infty}\in\left[-\infty,+\infty\right]$ as $n\to\infty$. Then,
$\lim_{n\to\infty}{\mathbb{P}}\left({\mathcal{G}}^{(k)}(n,m,p)\text{ \emph{is connected}}\right)=\lim_{n\to\infty}{\mathbb{P}}\left(\min\deg{\mathcal{G}}^{(k)}(n,m,p)\geq 1\right)=\exp\left(-e^{-\alpha_{\infty}}\right),$
where we compute $\exp\left(-e^{-(-\infty)}\right)=0$ and $\exp\left(-e^{-(+\infty)}\right)=1$.
Now, we aim to find a sufficient condition for $N$ to have
$\sum_{b\in B_{D}}{\mathbb{P}}_{\mathrm{Po}(N-{\varepsilon})}\left({\mathcal{G}}_{b}^{(D,k)}\text{ is disconnected}\right)\leq\frac{\delta}{2}.$ (9)
Let us choose the sequence $\alpha_{n}=\ln n$ in Lemma˜F.14, which guarantees that
${\mathbb{P}}({\mathcal{G}}^{(k)}(n,m,p)\text{ is connected})\to 1\quad\text{ as }n\to\infty.$
From the definition of the limit of a sequence, for any choice of $\delta>0$, let us define $n_{0}(\delta)>e>0$ such that
${\mathbb{P}}({\mathcal{G}}^{(k)}(n,m,p)\text{ is disconnected})\leq\delta,\qquad\forall n>n_{0}(\delta).$
Then, choosing all $n_{b}=\left|V_{b}\right|$ to be greater than $n_{0}\!\left(\frac{\delta}{2\left|B_{D}\right|}\right)$, we yield the bound (9). Note that a choice of $p$ larger than the threshold in Eq.˜8 will never change that ${\mathbb{P}}({\mathcal{G}}^{(k)}(n,m,p)\text{ is connected})\to 1$ as $n\to\infty$. Thus, since we choose $\alpha_{n}=\ln n$, it suffices to have
$p=1-\exp\left(-\frac{N-{\varepsilon}}{\left|{\mathbb{X}}\right|}\right)\geq\left(\frac{k!\,(2\ln n_{b})}{n_{b}}\right)^{\frac{1}{2k}}\cdot\frac{1}{\sqrt{\left|{\mathcal{X}}_{3}\right|}}$
for each $n_{b}$ ($\forall b\in B_{D}$). Since $1-e^{-u}\geq(1-\frac{1}{e})u$ for $0<u<1$ and $\frac{\ln u}{u}$ is a decreasing function for large enough $u>e$, it suffices to have
$\left(1-\frac{1}{e}\right)\frac{N-{\varepsilon}}{\left|{\mathbb{X}}\right|}\geq\left(\frac{k!\,(2\ln\hat{n})}{\hat{n}}\right)^{\frac{1}{2k}}\cdot\frac{1}{\sqrt{\left|{\mathcal{X}}_{3}\right|}}$
for $\hat{n}:=\min_{b\in B_{D}}n_{b}$. Plugging in ${\varepsilon}=\sqrt{2N\ln\left(\frac{2}{\delta}\right)}$ and $\left|{\mathbb{X}}\right|=\left|{\mathcal{X}}_{1}\!\times\!{\mathcal{X}}_{2}\right|\left|{\mathcal{X}}_{3}\right|$, we have
$N-\sqrt{2N\ln\left(\frac{2}{\delta}\right)}\geq\left(\frac{e}{e-1}\right)\cdot\left(\frac{k!\,(2\ln\hat{n})}{\hat{n}}\right)^{\!\!\frac{1}{2k}}\!\cdot\left|{\mathcal{X}}_{1}\!\times\!{\mathcal{X}}_{2}\right|\sqrt{\left|{\mathcal{X}}_{3}\right|},$
which is satisfied by
$N\geq 4\cdot\max\left\{\left(\frac{e}{e-1}\right)\cdot\left(\frac{k!\,(2\ln\hat{n})}{\hat{n}}\right)^{\!\!\frac{1}{2k}}\!\cdot\left|{\mathcal{X}}_{1}\!\times\!{\mathcal{X}}_{2}\right|\!\sqrt{\left|{\mathcal{X}}_{3}\right|},\;\ln\left(\frac{2}{\delta}\right)\right\}.$
In conclusion, for large enough $\hat{n}>\max\left\{n_{0}\!\left(\frac{\delta}{2\left|B_{D}\right|}\right),e\right\}$, provided that the training dataset size $N$ satisfies the inequality above, we finally have
$\ {\mathbb{P}}_{N}\left(\exists b\in B_{D}\text{ such that }{\mathcal{G}}_{b}^{(D,k)}\text{ is disconnected}\right)$
$\leq\sum_{b\in B_{D}}{\mathbb{P}}_{\mathrm{Po}(N-{\varepsilon})}\left({\mathcal{G}}_{b}^{(D,k)}\text{ is disconnected}\right)+\exp\left(-\frac{{\varepsilon}^{2}}{2N}\right)$
$\leq\frac{\delta}{2}+\frac{\delta}{2}=\delta.$
This concludes the proof of the theorem. ∎
Proof of Theorem˜F.8.
Even for $k=1$, we follow almost the same proof as that of Theorem˜F.7, except for the last few steps. Namely, we again observe that every ${\mathcal{G}}_{b}^{(D,1)}$ is a binomial random $1$-intersection graph ${\mathcal{G}}^{(1)}(n_{b},m,p)$ with parameters $n_{b}=\left|V_{b}\right|$, $m=\left|{\mathcal{X}}_{3}\right|$, and $p=1-\exp(-(N-{\varepsilon})/\left|{\mathbb{X}}\right|)$ (see the definition of binomial random intersection graph in Sec.˜F.3.2). To guarantee the union bound in Eq.˜9, however, we use a different lemma: Lemma˜F.15 (instead of Lemma˜F.14). In particular, we will use only the $\rho\leq 1$” part (which tightly matches our assumption $m=O(n_{b})$ ($b\in B_{D}$)) of the lemma below:
Lemma F.15 (singer1995random, Propositions 3.1–2, Theorem 3.3).
Let $k=1$. Suppose that $m=n^{\rho}$ for $\rho>0$ and
$p=\begin{dcases}\frac{\ln n+\alpha_{n}}{m}&\text{for }\rho\leq 1;\\ \sqrt{\frac{\ln n+\alpha_{n}}{mn}}&\text{for }\rho>1,\end{dcases}$
for any sequence $\{\alpha_{n}\}$ which attains a limit $\alpha_{\infty}\in\left\{-\infty,+\infty\right\}$ as $n\to\infty$. Then,
@@
Again, let us choose the sequence $\alpha_{n}=\ln n$ in Lemma˜F.15, which guarantees that ${\mathbb{P}}({\mathcal{G}}^{(1)}(n,m,p)\text{ is connected})\to 1$ as $n\to\infty$. From the definition of the limit of a sequence, for any choice of $\delta>0$, let us define $n_{0}(\delta)>1$ such that
${\mathbb{P}}({\mathcal{G}}^{(1)}(n,m,p)\text{ is disconnected})\leq\delta,\qquad\forall n>n_{0}(\delta).$
Then, choosing all $n_{b}=\left|V_{b}\right|$ (and thus $\hat{n}=\min_{b\in B_{D}}n_{b}$) to be greater than $n_{0}\!\left(\frac{\delta}{2\left|B_{D}\right|}\right)$, we yield the same bound (9). Now, by our choice $\alpha_{n}=\ln n$, it suffices to have
$p=1-\exp\left(-\frac{N-{\varepsilon}}{\left|{\mathbb{X}}\right|}\right)\geq\frac{2\ln n_{b}}{|{\mathcal{X}}_{3}|}$
for each $n_{b}$ ($\forall b\in B_{D}$). Since $1-e^{-u}\geq(1-\frac{1}{e})u$ for $0<u<1$ and $\ln u$ is an increasing function for $u>1$, it suffices to have
$\left(1-\frac{1}{e}\right)\frac{N-{\varepsilon}}{\left|{\mathbb{X}}\right|}\geq\frac{2\ln\check{n}}{|{\mathcal{X}}_{3}|}$
for $\check{n}:=\max_{b\in B_{D}}n_{b}$. Plugging in ${\varepsilon}=\sqrt{2N\ln\left(\frac{2}{\delta}\right)}$ and $\left|{\mathbb{X}}\right|=\left|{\mathcal{X}}_{1}\!\times\!{\mathcal{X}}_{2}\right|\left|{\mathcal{X}}_{3}\right|$, we have
$N-\sqrt{2N\ln\left(\frac{2}{\delta}\right)}\cdot\geq\left(\frac{e}{e-1}\right)(2\ln\check{n})\cdot\left|{\mathcal{X}}_{1}\!\times\!{\mathcal{X}}_{2}\right|,$
which is satisfied by
$N\geq 4\cdot\max\left\{\left(\frac{e}{e-1}\right)\cdot(2\ln\check{n})\!\cdot\!\left|{\mathcal{X}}_{1}\!\times\!{\mathcal{X}}_{2}\right|,\;\ln\left(\frac{2}{\delta}\right)\right\}.$
In conclusion, for large enough $\check{n}\geq\hat{n}>\max\left\{n_{0}\!\left(\frac{\delta}{2\left|B_{D}\right|}\right),1\right\}$, provided that the training dataset size $N$ satisfies the inequality above, we finally have
$\ {\mathbb{P}}_{N}\left(\exists b\in B_{D}\text{ such that }{\mathcal{G}}_{b}^{(D,1)}\text{ is disconnected}\right)$
$\leq\sum_{b\in B_{D}}{\mathbb{P}}_{\mathrm{Po}(N-{\varepsilon})}\left({\mathcal{G}}_{b}^{(D,1)}\text{ is disconnected}\right)+\exp\left(-\frac{{\varepsilon}^{2}}{2N}\right)$
$\leq\frac{\delta}{2}+\frac{\delta}{2}=\delta.$
This concludes the proof. ∎
Subsequently, we argue that the sample complexity upper bound obtained in Theorem˜F.7 (for $k\geq 2$) is nearly tight. [^7] We often say a complexity upper bound is tight if we can find a worst-case example whose complexity lower bound is almost identical to the upper bound. To this end, we show that the learner (we assumed in Sec.˜F.1) cannot avoid an incomplete coverage (i.e., $\overline{D}\not\subset\mathrm{Cover}_{k}(D)$) with high probability, with a dataset slightly smaller than our upper bound for certain instances of the 2-Hop task. In particular, it is enough to consider a subclass of the 2-Hop task satisfying a mild condition (Assumption˜F.6) that any two distinct intermediate states in $B$ never share the same output for more than $k_{\star}$(<k<k) tokens in ${\mathcal{X}}_{3}$. The formal statement of the tightness result is shown below:
Theorem F.16 (Near-Tightness of the Sample Complexity Upper Bound for $k\geq 2$, $\downarrow$).